OpenStudy (anonymous):

A rectangle has an area of 72 square units. The width of the rectangle is 9 units. The length of the rectangle is 2x + 4. What is the rectangle's length?

OpenStudy (anonymous):

a= l*w 72 = (2x+4) 9 72 = 18x + 36 72-36 = 36 36/18 = 2 x = 2

OpenStudy (anonymous):

length = 8